A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?

A. -3x + 4y = 3

B. -1.5x − 3.5y = -31.5

C. 2x + y = 20

D. -2.25x + y = -9.75


1
Expert's answer
2021-06-24T17:45:49-0400

the given line is

"-7x+3y=-21.5\\

3y=7x-21.5\\

y=\frac{7}{3}x-\frac{21.5}{3}"


here coefficient of x is 7/3 so the slope of the given line is

"m_1=\frac{7}{3}"


now the line PQ should be either parallel or perpendicular 

let the slope of the line PQ is  m2

so when m1=m2  then both lines are parallel

and when m1.m2=-1  then both lines are perpendicular 


now check the first option 

"-3x+4y=3\\

4y=3x+3\\

y=\frac{3}{4}x+\frac{3}{4}"


here coefficient of x is 3/4 so the slope of the first line is

"m_2= \frac{3}{4} \\

so \space here \space m_1\ne m_2 



."

so both lines are not parallel


now find its product

"m_1 \cdot m_2= \frac{7}{3}\cdot \frac{3}{4}\\

m_1 \cdot m_2= \frac{7}{4}\\

m_1 \cdot m_2 \ne -1\\"

so both lines are not perpendicular 


now check the second option 

"-1.5x-3.5y=-31.5"

"-3.5y=1.5x-31.5"

"3.5y=-1.5x+31.5"

"y=-\frac{1.5}{3.5}x+\frac{31.5}{3.5}"


"y=-\frac{15}{35}x+\frac{315}{35}"


"y=-\frac{3}{7}x+9"


"so\space here \space m_1\ne m_2"


so both lines are not parallel

"now\space find \space its \space product\\

m_1 \cdot m_2= \frac{7}{3}\left(-\frac{3}{7}\right)

{\color{Red} m_1 \cdot m_2= -1}"

therefore both lines are perpendicular 


so option B is correct

"{\color{Red} -1.5x-3.5y=-31.5}

."